OpenStudy (anonymous):

evaluate the limit x-->0 |x|/x^3

OpenStudy (anonymous):

which direction ?

OpenStudy (anonymous):

sorry limit when x approaches 0 from the left

OpenStudy (anonymous):

maybe it doesn't matter since the limit doesn't exist in any case

OpenStudy (anonymous):

from the left you get minus infinity , since if x<0 |x|=-x and you get \[\frac{-x}{x^3}=-\frac{1}{x^2}\] limit is minus infinity